Adjapong is 3times as old as atsu. Three
Years ago he was four times as old as
Atsu. How old is each boy now?

Found 2 solutions by ikleyn, solver91311:
Answer by ikleyn(52790)

.

If the younger is x years old, then the older is 3x years old now.


Three years ago, the younger was x-3 years old, while the older was 3x-3 years old.


From the condition

   3x - 3 = 4*(x-3).


Simplify and solve

    3x - 3 = 4x - 12

    12 - 3 = 4x - 3x

     9     = x


ANSWER.  Atsu is 9 years old;  Adjapong is 9*3 = 27 years old.

Solved.

----------------

There is a bunch of lessons on age word problems in this site:
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - One unusual age word problem
    - Miscellaneous age problems

Read them attentively and become an expert in this field.




Answer by solver91311(24713)



Let the older one's age be represented by and the younger one's age be represented by .





Solve for and


John

My calculator said it, I believe it, that settles it